OpenStudy (anonymous):

3 divided by square root of 5, step by step please thanks

OpenStudy (anonymous):

mimi (:

OpenStudy (mimi_x3):

ohh um , hello(:

OpenStudy (anonymous):

lol idk how to put up the square root symbol lol

OpenStudy (mimi_x3):

lols , go to equation icon lmao

OpenStudy (anonymous):

hey angela (:

OpenStudy (angela210793):

3/sqr5 multiply up and down with aqr5 3sqr5/(sqr5*sqr5)=3sqr5/5 Hi Fresh! :D

OpenStudy (anonymous):

hey do you know how to show the symbols its more simple that way lol

OpenStudy (anonymous):

its confusing like this lol

OpenStudy (mimi_x3):

\[3/\sqrt{5} \times \sqrt{5}/\sqrt{5}\] = \[3\sqrt{5}/ 5\]

OpenStudy (mimi_x3):

lols , it took me a long time to type it up xDD

OpenStudy (anonymous):

3/sqr5 =3*sqr5/sqr5*sqr5 =3sqr5